# Random Schreier graphs as expanders

### v6

@@ A.1 Proofs of Theorems 1 and 3
Let $q$ be a prime, $k$ a natural integer and $G=Sch(GL_{k}(\mathbb{F}_{q})\circlearrowleft(\mathbb{F}_{q}^{k})^{*},S)$ be a Schreier graph with $S$ a random subset of $d$ elements of $GL_{k}(\mathbb{F}_{q})$ and their inverses. Hence, we obtain a $2d$ regular graph whose size is $n=q^{k}-1$. By mapping every element of $(\mathbb{F}_{q}^{k})^{*}$ to an element of $[\![1,\ q^{k}-1]\!]$ (e.g. with $f:(x_{1},\dots,x_{k})\mapsto\sum_{i=0}^{k-1}x_{i}q^{i}$), we can associate every pair of vertices with a coordinate of a matrix. This way, we can define $M$, the normalised adjacency matrix of $G$. Let
@@
Consider a random walk of $2m$ steps starting at vertex $i$. Then the $(i,i)$ coordinate of $M^{2m}$ corresponds to the number of closed walks starting at vertex $i$ of size $2m$ divided by $(2d)^{2m}$, since $(2d)^{2m}$ is the number of paths of size $2m$ starting at $i$. Therefore, this is the probability (denoted $P_{ii}$) of returning to the vertex $i$ after $2m$ steps of the random walk. Since $Trace(M^{2m})$ is equal to the sum of all of these quantities and since the expected $(i,i)$ coordinate of $M^{2m}$ is the same for every $i$ we get
$\mathbf{E}(Trace(M^{2m}))=n\mathbf{E}(P_{11}).$
@@
$\sum_{i=1}^{n}\mu_{i}^{2m}=Trace(M^{2m}).$
Thus, since $\mu_{i}^{2m}\geq 0$ and $\mu_{1}=1$, we get $|\mu_{2}|\leq(Trace(M^{2m})-1)^{1/2m}$, which implies by Jensen’s inequality
@@
Given the starting vertex (say $v_{1}=f^{-1}(1)$), the random walk can be seen as the product of $2m$ elements of $S$ chosen uniformly and independently at random. We denote this product $\omega=w_{2m}.w_{2m-1}\dots w_{2}.w_{1}$ and each of its matrices represents the choice of a particular neighbour for all vertices. Hence, if the first vertex on the path is $v_{1}$, the second will be $v_{2}=s_{i_{1}}.v_{1}$ (where $s_{i_{1}}$ is the value of $w_{1}$), the third $v_{3}=s_{i_{2}}.v_{2}$ (where $s_{i_{2}}$ is the value of $w_{2}$) and so on; the last vertex of the path is then $\omega.v_{1}$.
@@
We reuse here the main conceptual idea of the proof in [6]. The value of $\omega.v_{1}$ depends on two types of random choices: on the random choice of the word $\omega=w_{2m}.w_{2m-1}\dots w_{2}.w_{1}$ where each letter is chosen at random in $\{s_{1},s^{-1}_{1},\dots,s_{d},s_{d}^{-1}\}$, and the random choice of a matrix in $GL_{k}(\mathbb{F}_{q})$ for each $s_{i}$. These two choices are independent. We may sample at first the words $\omega$ and only then choose the matrices $s_{j}$. We prefer not to sample the entire value of each $s_{j}$ in “one shot” but reveal the values of these matrices (better to say, the values of the linear operators corresponding to these matrices) little by little, as it is needed. Thus, starting at vertex $v_{1}$, instead of choosing at random in $GL_{k}(\mathbb{F}_{q})$ the entire matrix $w_{1}=s_{i_{1}}$, we only determine the result of the product $v_{2}=s_{i_{1}}.v_{1}$. This choice does not determine completely the matrix $s_{i}$ but imposes a linear constraint on the matrix elements of $s_{i_{1}}$. The same letter $w_{1}$ may appear in the word $\omega$ several times. Each time the same letter $w_{1}$ appears in the word $\omega$ and, therefore, the matrix $s_{i_{1}}$ is encountered on the path, we must define the action of this matrix on some new vector $x$. We choose the result of $s_{i_{1}}.x$ by extending the partial definition of $s_{i_{1}}$, which means an extension of the linear constraints on $s_{i}$ fixed earlier. In a similar way, we define step by step the other matrices $s_{j}$ that are involved in $\omega$. We need to understand the distribution of the vector $v_{2m}=\omega.v_{1}$ that we obtain at the end of this procedure (and the probability of the event $v_{2m}=v_{1}$). In the next paragraphs we analyse this distribution.
@@
Consider a matrix $s\in S$ that has been already encountered on the path refined by $\omega$, and we have already defined the action of $s$ on $t$ different vertices. Assume that we encounter the same matrix $s$ once again, and we must define the product $s.x$ for some one more vector $x\in(\mathbb{F}_{q}^{k})^{*}$). In the permutation model, as stated in [6] this would be a uniform distribution over the $n-t$ vertices that have not been earlier assigned to the partially defined permutation $s$. However, in our construction, even if $x$ is totally new to $s$, the result of $s.x$ may not be necessarily undetermined. Indeed, if $x$ is linearly dependent from the vectors that we have already met, we would have
@@
would be completely determined by our previous random choices, and would not give any new information about $s$. Intuitively, we would say that the step that leads from $x$ to $s.x$ is not free. In order characterise formally what it means for a step to be free, we need to introduce the following set: let $s$ be a matrix of $S$, $\omega=w_{2m}\dots w_{1}$, $v_{1}$ the starting vertex and $v_{j+1}=w_{j}v_{j}$. Then we define
@@
Definition A.1 (free and forced step).
@@
Alternatively, instead of saying that a step $i$ is free, we will say that the vector obtained after this step is free (namely the $i+1$-th vector, $w_{i}.v_{i}$). The following lemma justifies this terminology, and will be used systematically in the rest of the paper.
Lemma A.1.1.
Let $s=w_{i}$ for a step $i$ and $t$ be the dimension of $\Sigma_{s}(i)$. Then, if $v_{i}\notin\Sigma_{s},\ v_{i+1}=s.v_{i}$ can be chosen uniformly at random among the $q^{k}-q^{t}$ vectors that do not belong to $s.\Sigma_{s}(i)$.
Proof.
The choice of a non degenerate matrix $s$ of size $k\times k$ is the same as the choice of an bijective linear operator from $\mathbb{F}_{q}^{k}$ to $\mathbb{F}_{q}^{k}$. To specify a linear operator, we only need to define it on vectors of any basis in $\mathbb{F}_{q}^{k}$. Let $x_{1},\dots,x_{t}$ be a basis of $\Sigma_{s}(i)$. By the assumption of the lemma, vector $v_{i}$ is linearly independent with $x_{1},\dots,x_{t}$. Therefore, we can let $x_{t+1}=v_{i}$ and then extend $x_{1},\dots,x_{t},x_{t+1}$ to a basis in the space $\mathbb{F}_{q}^{k}$ with some $x_{t+2},\ldots,x_{k}$.
To define $s$, we should specify one by one linearly independent vectors $y_{1}=s.x_{1},y_{2}=s.x_{2},\ldots,y_{k}=s.x_{k}$. We have $q^{k}-1$ possibilities to choose $y_{1}$ (any non zero vector), $q^{k}-q$ possibilities to choose $y_{2}$ (any vector linearly independent with the fixed $y_{1}$), $q^{k}-q^{2}$ possibilities to choose $y_{3}$ (any vector linearly independent with $y_{1}$ and $y_{2}$), and so on. In particular, if we have fixed the values $y_{i}=s.x_{i}$ for $i=1,\ldots,t$, then it remains $q^{k}-q^{t}$ available options to choose $y_{t+1}$ (which is the same as $v_{i+1}$ in our notation).
@@
This implies a sort of transitivity of the group action which is stronger than the simple transitivity, but weaker than the $k$-transitivity: for all $t\leq k$, if $(x_{1},\dots,x_{t})$ and $(y_{1},\dots,y_{t})$ are two families of linearly independent elements of $\mathbb{F}_{q}^{k}$ then, there exists an element $s$ of $GL_{k}(\mathbb{F}_{q})$ such that for all $i\leq t$, $s.x_{i}=y_{i}$.
As a direct consequence of this lemma, we can formulate the following corollary:
Corollary 1.
If a matrix $s$ appears only once in $\omega$ at step $i$, $s.v_{i}=v_{i+1}$ is chosen uniformly at random among every vertex of the graph.
Proof.
Indeed, when this happens, we can rewrite $\omega$ as $AsB$ with $A$ and $B$ some invertible matrices. Then we can expose the coefficients of $A$ and $B$. The key point is that those matrices are independent from $s$, which means that we still have no information about $s$. This way, since $v_{i}=B.v_{1}$, we have $\mathbf{P}(\omega.v_{1}=v_{1})=\mathbf{P}(s.v_{i}=A^{-1}v_{1})=\frac{1}{n}$. ∎
The probability we have just found is conditioned by the structure of $\omega$, hence it needs to be multiplied by the probability of the condition. This is only dependent on the properties of $\omega$ which can be seen as an element of a free group generated by the set $\{s_{1},\dots s_{d}\}$. Let us remind that $\omega$ can be seen as a words whose letters are taken from $\{s_{1},s^{-1}_{1},\dots s_{d},s^{-1}_{d}\}$. In order to estimate the total probability of having a closed walk, we subdivide the space of such words in a few events whose probabilities will be determined. Those events are chosen so the conditional probability is more convenient to estimate. This will be done later. Let us define our events:
  • $X_{1}$ : “at least one letter appears exactly once”
@@
These events form a partition of the set of all possible words, whose size is $(2d)^{2m}$. In order to finish the proof of Theorem 3, let us leave aside all the events but $X_{1}$ and consider its complementary. We bound their probability using an argument similar to that of [3]. We observe that a word that belongs to $\overline{X_{1}}$ has at most $m$ different letters in it, hence we have at most $\binom{d}{m}$ ways of choosing those letters in the alphabet. When choosing each letter at random, the probability that all of them are in the right set is $(\frac{m}{d})^{2m}$. Hence
The probability we are looking for is then
$\leq\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})+\mathbf{P}(\overline{X_{1}})\leq\frac{1}{n}+\bigg{(}e\frac{m}{d}\bigg{)}^{m}.$
We set $m=\ln n$ to minimise the bound. Substituting the above quantity back into equation 3 completes the proof of Theorem 3.
In order to find a tighter bound, we use a more careful analysis. We consider again all of our events. We are going to represent the probability $P(\omega.v_{1}=v_{1})$ as the sum
@@
For $i=1,2,3$ and 4 we estimate separately $P(X_{i})$ and $P(\omega.v_{1}=v_{1}\mid\omega\in X_{i})$, and we estimate the value of the product $P(\omega.v_{1}=v_{1}\mid\omega\in X_{2}^{\prime})\cdot P(X_{2}^{\prime})$ as a whole. The sum of these bounds result in the proof of Theorem 1.
@@
We have already explained that $\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})=\frac{1}{n}$. It remains to bound this probability conditioned to the other events. In what is next, we will set $q=2$. Taking a bigger field might weaken the bounds and complicate the analysis. We start with $X_{2}$.
Lemma A.1.2.
@@
By taking the maximum of all these bounds, we conclude the proof. ∎
@@
Now we suppose $x\not=y$. Then if $y^{\prime}=x^{\prime}$ we have $y^{\prime\prime}=x\neq y$. If $y^{\prime}\neq x^{\prime}$ (which happens with probability $\frac{n-1}{n}$, $y^{\prime\prime}$ is free, and its value is uniformly distributed among the $n-1$ remaining vectors. Therefore, when we have this configuration of random matrices in $\omega$, the probability of having a closed walk is $\frac{1}{n}$. ∎
@@
If the $i$ pairs of letters in $\omega$ do not contain the pattern (6) but one of other letters appear in $\omega$ exactly thee times, then the probability to have a closed path is at most $\frac{5}{n}$, as shown in Lemma A.1.3. This case contributes to the resulting probability at most
@@
The sum of these expressions is larger than $P_{11}$ defined at the beginning of this section. By replacing it in equation 3, we complete the proof. It is easy to see that the rough bounding used in the proof of Theorem 3 gives a larger bound than that of Theorem 2.
A.2 Proof of Theorem 2
@@
$22\mu_{3}^{2m}\leq\sum_{i=1}^{n-1}2\mu_{2i+1}^{2m}=Trace(M^{2m})-$ 2 .
As we have seen in section 3.1, the expected value of $Trace(M^{2m})$ is the sum of the probability of getting of closed path of size $2m$, starting on each vertex. We note this probability $P_{ii}$ for the vertex $i$. It is the same for every vertex, hence we get, by using Jensen’s inequality
$\mathbf{E}(|\mu_{3}|)\leq\bigg{(}\frac{1}{2}(\mathbf{E}(Trace(M^{2m}))-2)\bigg{)}^{\frac{1}{2m}}=(nP_{11}-1)^{\frac{1}{2m}}.$ (7)
@@
We first explain why the construction for even degree regular bipartite graphs gives the same bound as Theorem 1. Here, a random walk can be represented as a sequence of matrices of $D\cup D^{-1}$. This is because every vertex $x$ is connected to $s.x$ and $s^{-1}.x$. Each element of the sequence is chosen independently of the others. It is then easy to see that the structure of the walk is exactly the same as in the non bipartite case: a uniformly random sequence of $2m$ matrices from $D\cup D^{-1}$. The same proof can then be applied to this sequence, the elements of the sequence will then behave the same way as in the preceding section.
@@
  • $Y_{1}$ : “at least one letter appears exactly once”
  • $Y_{2}$ : $\overline{Y_{1}}$wedge ”at least one letter appears exactly twice”
  • $Y_{3}$ : “no letter appear once or twice, at least one letter appears exactly three times”
  • $Y_{4}$ : “no letter appear once, twice, nor three times”
Up to this point, it is not hard to understand why the bound of Theorem 3 holds. Indeed, the sign of the matrices do not play any role in the proof, so the probability of $\overline{Y_{1}}$ can be bounded by the same quantity as in the previous section. In addition, the probability of having a closed walk conditioned to $Y_{1}$ is also $\frac{1}{n}$ ($n$ is the size of a partition). Therefore, Theorem 3 applies to bipartite regular graphs.
@@
We have already shown that when one letter appears exactly once, the probability of having a closed walk is $\frac{1}{n}$. Similarly, when a letter appears exactly three times, the probability of getting a closed walk is less than $\frac{5}{n}$. Indeed, in the proof of Lemma A.1.3, all possible configurations of signs for the letter that appears three times are considered (e.g. $\omega=\dots s\dots s^{-1}\dots s\dots$ or $\dots s\dots s\dots s^{-1}\dots$). No assumption is done on their respective probabilities to occur. These probabilities may or may not be different in the bipartite setting. Since this bound ($\frac{5}{n}$) is the maximum over all the probabilities of getting a closed walk with each configuration of signs, the resulting bound for the probability of getting a closed walk in the bipartite case remains the same. Hence we get

### v7

@@ 0.B.1 Proof of Proposition 1
Let H be a group acting tansitively on a set $V$ and let $G$ the Schreier graph of $S\subset H$ acting on $V$ with $S$ a random multiset of $d$ elements. Hence, we obtain an undirected $2d$ regular graph whose size is $n=|V|$. By mapping every element of $V$ to an element of $[\![1,\ n]\!]$ (with some bijective function $f$), we can associate every pair of vertices with a coordinate of a matrix. This way, we can define $M$, the normalised adjacency matrix of $G$. Let
@@
Consider a random walk of $2m$ steps starting at vertex $i$. Then the $(i,i)$ coordinate of $M^{2m}$ corresponds to the number of closed walks starting at vertex $i$ of size $2m$ divided by $(2d)^{2m}$, since $(2d)^{2m}$ is the number of paths of size $2m$ starting at $i$. Therefore, this is the probability (denoted $P_{ii}$) of returning to the vertex $i$ after $2m$ steps of the random walk. Since $\textit{Trace}(M^{2m})$ is equal to the sum of all of these quantities and since the expected $(i,i)$ coordinate of $M^{2m}$ is the same for every $i$ we get
$\mathbf{E}(\textit{Trace}(M^{2m}))=n\mathbf{E}(P_{11}).$
@@
$\sum_{i=1}^{n}\mu_{i}^{2m}=\textit{Trace}(M^{2m}).$
Thus, since $\mu_{i}^{2m}\geq 0$ and $\mu_{1}=1$, we get $|\mu_{2}|\leq(\textit{Trace}(M^{2m})-1)^{1/2m}$, which implies by Jensen’s inequality
@@
Given the starting vertex (say $v_{1}=f^{-1}(1)$), the random walk can be seen as the product of $2m$ group elements that belong to $S$ or their inverses (that is, $S\cup S^{-1}$) chosen uniformly and independently at random. We denote this product of group elements $\omega=w_{2m}.w_{2m-1}\dots w_{2}.w_{1}$ and each $w_{i}$ represents the choice of a particular neighbour for all vertices. Hence, if the first vertex on the path is $v_{1}$, the second will be $v_{2}=s_{i_{1}}.v_{1}$ (where $s_{i_{1}}$ is the value of $w_{1}$), the third $v_{3}=s_{i_{2}}.v_{2}$ (where $s_{i_{2}}$ is the value of $w_{2}$) and so on; the last vertex of the path is then $\omega.v_{1}$.
In what follows we call by literals the $2d$ elements of the set
$\{s_{1},s^{-1}_{1},\dots,s_{d},s_{d}^{-1}\}$
where each letter $s_{j}$ (for $j=1,\ldots,d$) appears in two ways: as the literal $s_{j}$ and as the literal $s_{j}^{-1}$. Thus, a product of elements of $S\cup S^{-1}$ $\omega=w_{2m}w_{2m-1}\dots w_{2}w_{1}$ can be identified with a sequence of $2m$ literals.
We reuse here the main conceptual idea of the proof in [6]. The value of $\omega.v_{1}$ depends on two types of random choices: on the random choice of the word $\omega=w_{2m}.w_{2m-1}\dots w_{2}.w_{1}$ where each literal is chosen at random in $\{s_{1},s^{-1}_{1},\dots,s_{d},s_{d}^{-1}\}$, and the random choice of an element in $H$ for each $s_{i}$. These two choices are independent. It is usefull here to sample at first the words $\omega$ and only then choose the elements $s_{j}$.
We need the following simple lemma in order to estimate the probability of a closed walk. It applies to any transitive group action; this statement also applies if $s_{j}$ are chosen among Toeplitz matrices (even though the set of all non-singular Toeplitz matrices with matrix multiplication is not a group).
Lemma 1
Let $s$ be an element of a group acting transitively on a set $V$. If letter $s$ appears in a word $\omega$ only once, at some position $i$, then $v_{i+1}=s.v_{i}$ is chosen uniformly at random among all elements of $V$. Moreover, for such a word $\omega$, the probability of the event $\omega.v_{1}=v_{1}$ is equal to $1/n$.
The same property is true if $s$ is chosen as a random Toeplitz matrix acting on the set $V$ of non-zero vectors.
Proof
Indeed, if letter $s$ appears in $\omega$ exactly once, we can rewrite $\omega$ as $AsB$, where $A$ and $B$ are elements of the group chosen chosen at random (not necessary uniformly and independently of each other). The key observation is that $A$ and $B$ are independent from $s$. So we can at first choose the values of $A$ and $B$. At this stage we still have no information about $s$.
Further, since the action on $V$ is transitive, for every $v_{i}$ and a randomly chosen $s$, the values $s.v_{i}$ are uniformly distributed on $V$. Thus, as we have fixed $v_{i}=B.v_{1}$, we have $\mathbf{P}(\omega.v_{1}=v_{1})=\mathbf{P}(s.v_{i}=A^{-1}v_{1})=\frac{1}{n}$.
The same is true if $s$ is a Toeplitz matrix. Indeed, a Toeplitz matrix has at least one completely free column. Hence the result of the product $s.v_{i}$ is uniformly distributed (see Lemma 2 in the next section).
The probability that we have just found is computed for a fixed word $\omega$ under the condition that this word contains a letter that appears there exactly once. We denote this condition by $X_{1}$ (it is an event on $\omega$’s probability space). It remains to compute the fraction of words having this property. Let us remind that $\omega$ can be seen as a word whose symbols (literals) are taken at random from $\{s_{1},s^{-1}_{1},\dots s_{d},s^{-1}_{d}\}$.
In order to finish the proof of Proposition 1, we bound the number of words in which no letter appears exactly once using an argument similar to that of [3]. We observe that for a word that belongs to $\overline{X_{1}}$ there are at most $m$ different indices $j$ such that $s_{j}$ or $s_{j}^{-1}$ (or both) appear in the word. Hence, we have at most $\binom{d}{m}$ ways of choosing those letters in the alphabet. When choosing each letter at random, the probability that all of them are in the right set is $(\frac{m}{d})^{2m}$. Thus,
The probability we are looking for is then
$\leq\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})\cdot 1+1\cdot\mathbf{P}(\overline{X_{1}})\leq\frac{1}{n}+\bigg{(}e\frac{m}{d}\bigg{)}^{m}.$
We set $m=\ln n$ to minimise the bound. Substituting the above quantity back into equation (3) completes the proof of Proposition 1.
0.B.2 Proof of Theorem 0.A.1
We now turn to proove Theorem 0.A.1. Let $k$ be a natural integer and $G$ the Schreier graph of $S\subset\textit{GL}_{k}(\mathbb{F}_{2})$ acting on $(\mathbb{F}_{2}^{k})^{*}$ by matrix-vector product, with $S$ a random multiset of $d$ elements. $G$ is thus a $2d$-regular graph of $12^{k}-$ 1 vertices. Let $M$ be the normalized adjacency matrix of $G$.
We reuse the ideas from the preceding section. Let $\omega=w_{2m}\dots w_{1}$ be a product of matrices chosen at random from $S\cup S^{-1}$. As before, $\omega$ represents a random walk in $G$. Here again, the choice of $S$ and the choice of $\omega$ are independent. Hence we can sample at first the words $\omega$ and only then choose the matrices that are in $S$.
@@
Following the approach in [6], we prefer not to sample the entire value of each $s_{j}$ in “one shot” but reveal the values of these matrices (better to say, the values of the linear operators corresponding to these matrices) little by little, as it is needed. Thus, starting at vertex $v_{1}$, instead of choosing at random in $\textit{GL}_{k}(\mathbb{F}_{2})$ the entire matrix $w_{1}=s_{i_{1}}$, we only determine the result of the product $v_{2}=s_{i_{1}}.v_{1}$. This choice does not determine completely the matrix $s_{i}$ but imposes a linear constraint on the matrix elements of $s_{i_{1}}$. The same letter $w_{1}$ may appear in the word $\omega$ several times. Each time the same letter $w_{1}$ appears in the word $\omega$ and, therefore, the matrix $s_{i_{1}}$ is encountered on the path, we must define the action of this matrix on some new vector $x$. We choose the result of $s_{i_{1}}.x$ by extending the partial definition of $s_{i_{1}}$, which means an extension of the linear constraints on $s_{i}$ fixed earlier. In a similar way, we define step by step the other matrices $s_{j}$ that are involved in $\omega$. We need to understand the distribution of the vector $v_{2m}=\omega.v_{1}$ that we obtain at the end of this procedure (and the probability of the event $v_{2m}=v_{1}$). In the next paragraphs we analyse this distribution. This will lead to a stronger version of Lemma 1 in the case of Schreier graphs of $\textit{GL}_{k}(\mathbb{F}_{2})$.
@@
Consider a matrix $s\in S$ that has been already encountered on the path defined by $\omega$, and we have already defined the action of $s$ on $t$ different vertices. Assume that we encounter the same matrix $s$ once again, and we must define the product $s.x$ for some one more vector $x\in(\mathbb{F}_{2}^{k})^{*}$). In the permutation model, as stated in [6] this would be a uniform distribution over the $n-t$ vertices that have not been earlier assigned to the partially defined permutation $s$. However, in our construction, even if $x$ is totally new to $s$, the result of $s.x$ may not be necessarily undetermined. Indeed, if $x$ is linearly dependent from the vectors that we have already met, we would have
@@
would be completely determined by our previous random choices, and would not give any new information about $s$. Intuitively, we would say that the step that leads from $x$ to $s.x$ is not free. In order to characterise formally what it means for a step to be free, we need to introduce the following set: let $s$ be a matrix of $S$, $\omega=w_{2m}\dots w_{1}$, $v_{1}$ the starting vertex and $v_{j+1}=w_{j}v_{j}$. Then we define
@@
Definition 1 (free and forced step)
@@
Alternatively, instead of saying that a step $i$ is free, we will say that the vector obtained after this step is free (namely the $(i+1)$-th vector, $w_{i}.v_{i}$). The following lemma justifies this terminology, and will be used systematically in the rest of the paper. We prove this lemma for invertible matrices over finite fields of prime size $q$. However, our result uses this only for fields of size $2$.
Lemma 2
Let $s=w_{i}$ for a step $i$ and $t$ be the dimension of $\Sigma_{s}(i)$. Then, if $v_{i}\notin\Sigma_{s},$ then $v_{i+1}=s.v_{i}$ can be chosen uniformly at random among the $q^{k}-q^{t}$ vectors that do not belong to $s.\Sigma_{s}(i)$.
Proof
The choice of a non degenerate matrix $s$ of size $k\times k$ is the same as the choice of a bijective linear operator from $\mathbb{F}_{q}^{k}$ to $\mathbb{F}_{q}^{k}$. To specify a linear operator, we only need to define it on vectors of any basis in $\mathbb{F}_{q}^{k}$. Let $x_{1},\dots,x_{t}$ be a basis of $\Sigma_{s}(i)$. By the assumption of the lemma, vector $v_{i}$ is linearly independent with $x_{1},\dots,x_{t}$. Therefore, we can let $x_{t+1}=v_{i}$ and then extend $x_{1},\dots,x_{t},x_{t+1}$ to a basis in the space $\mathbb{F}_{q}^{k}$ with some $x_{t+2},\ldots,x_{k}$.
To define $s$, we should specify one by one linearly independent vectors $y_{1}=s.x_{1},y_{2}=s.x_{2},\ldots,y_{k}=s.x_{k}$. We have $q^{k}-1$ possibilities to choose $y_{1}$ (any non zero vector; it is also true for Toeplitz matrices since at least one column is free [??]), $q^{k}-q$ possibilities to choose $y_{2}$ (any vector linearly independent with the fixed $y_{1}$), $q^{k}-q^{2}$ possibilities to choose $y_{3}$ (any vector linearly independent with $y_{1}$ and $y_{2}$), and so on. In particular, if we have fixed the values $y_{i}=s.x_{i}$ for $i=1,\ldots,t$, then it remains $q^{k}-q^{t}$ available options to choose $y_{t+1}$ (which is the same as $v_{i+1}$ in our notation).
@@
This implies a sort of transitivity property of the group action which is stronger than the simple transitivity, but weaker than the $k$-transitivity: for all $t\leq k$, if $(x_{1},\dots,x_{t})$ and $(y_{1},\dots,y_{t})$ are two families of linearly independent elements of $\mathbb{F}_{q}^{k}$ then, there exists an element $s$ of $\textit{GL}_{k}(\mathbb{F}_{q})$ such that for all $i\leq t$, $s.x_{i}=y_{i}$.
In order to estimate the total probability of having a closed walk, we subdivide the space of such words in a few subsets (events) and then estimate probabilities of each of them. We chose these subsets so that it will be easier to estimate the conditional probability to get closed path, as we show later. Let us define our events:
  • $X_{1}$ : “at least one letter appears in the word exactly once”
@@
These events form a partition of the set of all possible words, whose size is $(2d)^{2m}$. In the proof of Proposition 1, we only subdivided the space in two parts: $X_{1}$ and $\overline{X_{1}}$. In order to get a tighter bound, we need a more careful analysis of the combinatorial structure of $\omega$ and the corresponding conditional probabilities. We consider again all of our five events. We are going to represent the probability $P(\omega.v_{1}=v_{1})$ as the sum
@@
For $i=1,2,3$ and 4 we estimate separately $P(X_{i})$ and $P(\omega.v_{1}=v_{1}\mid\omega\in X_{i})$, and we estimate the value of the product $P(\omega.v_{1}=v_{1}\mid\omega\in X_{2}^{\prime})\cdot P(X_{2}^{\prime})$ as a whole. The sum of these bounds result in the proof of Theorem 0.A.1.
@@
We have already explained that $\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})=\frac{1}{n}$ (see Lemma 1). It remains to bound this probability conditioned to the other events. We start with $X_{2}$.
Lemma 3
@@
By taking the maximum of all these bounds, we conclude the proof.
@@
Now we suppose $x\not=y$. Then if $y^{\prime}=x^{\prime}$ we have $y^{\prime\prime}=x\neq y$. If $y^{\prime}\neq x^{\prime}$ (which happens with probability $\frac{n-1}{n}$, $y^{\prime\prime}$ is free, and its value is uniformly distributed among the $n-1$ remaining vectors. Therefore, when we have this configuration of random matrices in $\omega$, the probability of having a closed walk is $\frac{1}{n}$.
@@
If the $i$ pairs of letters in $\omega$ do not contain the pattern (6) but one of other letters appear in $\omega$ exactly thee times, then the probability to have a closed path is at most $\frac{5}{n}$, as shown in Lemma 4. This case contributes to the resulting probability at most
@@
The sum of these expressions is larger than $P_{11}$ defined at the beginning of this section. By replacing it in equation 3, we complete the proof. It is easy to see that the rough bounding used in the proof of Proposition 1 gives a larger bound than that of Theorem 0.A.2.
0.B.3 Proof of Theorem 0.A.2
@@
$22\mu_{3}^{2m}\leq\sum_{i=1}^{n-1}2\mu_{2i+1}^{2m}=\textit{Trace}(M^{2m})-$ 2 .
As we have seen in section 3.1, the expected value of $\textit{Trace}(M^{2m})$ is the sum of the probability of getting of closed path of size $2m$, starting on each vertex. We note this probability $P_{ii}$ for the vertex $i$. It is the same for every vertex, hence we get, by using Jensen’s inequality
$\mathbf{E}(|\mu_{3}|)\leq\bigg{(}\frac{1}{2}(\mathbf{E}(\textit{Trace}(M^{2m}))-2)\bigg{)}^{\frac{1}{2m}}=(nP_{11}-1)^{\frac{1}{2m}}.$ (7)
@@
We first explain why the construction for even degree regular bipartite graphs gives the same bound as Theorem 0.A.1. Here, a random walk can be represented as a sequence of matrices of $D\cup D^{-1}$. This is because every vertex $x$ is connected to $s.x$ and $s^{-1}.x$. Each element of the sequence is chosen independently of the others. It is then easy to see that the structure of the walk is exactly the same as in the non bipartite case: a uniformly random sequence of $2m$ matrices from $D\cup D^{-1}$. The same proof can then be applied to this sequence, the elements of the sequence will then behave the same way as in the preceding section.
@@
  • $Y_{1}$ : “at least one letter appears exactly once"
  • $Y_{2}$ : $\overline{Y_{1}}$wedge "at least one letter appears exactly twice"
  • $Y_{3}$ : “no letter appear once or twice, at least one letter appears exactly three times"
  • $Y_{4}$ : “no letter appear once, twice, nor three times"
Up to this point, it is not hard to understand why the bound of Theorem 1 holds. Indeed, the sign of the matrices do not play any role in the proof, so the probability of $\overline{Y_{1}}$ can be bounded by the same quantity as in the previous section. In addition, the probability of having a closed walk conditioned to $Y_{1}$ is also $\frac{1}{n}$ ($n$ is the size of a partition). Therefore, Proposition 1 applies to bipartite regular graphs.
@@
We have already shown that when one letter appears exactly once, the probability of having a closed walk is $\frac{1}{n}$. Similarly, when a letter appears exactly three times, the probability of getting a closed walk is less than $\frac{5}{n}$. Indeed, in the proof of Lemma 4, all possible configurations of signs for the letter that appears three times are considered (e.g. $\omega=\dots s\dots s^{-1}\dots s\dots$ or $\dots s\dots s\dots s^{-1}\dots$). No assumption is done on their respective probabilities to occur. These probabilities may or may not be different in the bipartite setting. Since this bound ($\frac{5}{n}$) is the maximum over all the probabilities of getting a closed walk with each configuration of signs, the resulting bound for the probability of getting a closed walk in the bipartite case remains the same. Hence we get